# Scalaron dark matter dynamics: effects of Higgs non-minimal coupling to gravity

### v1

@@ 3 Scalaron dynamics : effects of Higgs interactions
With the general setup explained earlier, we are all set to discuss two scenarios : in one of which the scalaron dynamics is governed by the Higgs quartic self interaction ($\lambda$) and in the other one, the non-minimal coupling to gravity ($\xi$) decides the dynamics along with an independent initial condition for the scalaron field. We solve Eq.16 and Eq.17 to demonstrate different dynamics at play, leading the scalaron to be a viable DM candidate. Consequently, the scalaron mass required to saturate the observed dark matter relic density differs widely, depending on the underlying dynamics.
@@ 3.1 Scenario I : $3\xi m^{2}\neq\lambda v^{2}$
In this scenario, the scalaron stays at its natural minima (as in Eq.12) before the EWPT and only starts moving as $T\leq T_{c}$. The scalaron, initially coupled to the SM thermal bath via the trilinear coupling, eventually decouples. Then it oscillates coherently as a matter field at temperatures well below $T_{c}$. This is a sheer contrast to axion misalignment scenario, in which the oscillation takes place depending only on the ratio of the Hubble parameter and the axion mass. Here, the onset of the oscillation is governed by the scalaron-Higgs interaction in addition to the Hubble to mass ratio. We demonstrate salient features of scalaron dynamics by solving Eq.16 semi-analytically with another change of variable, i.e. $\chi=\tilde{\phi}-\tilde{T}^{2}$, which yields,
@@
The equation can be solved analytically in the regime $\tilde{T}\leq 1$, for a constant $g_{*}(\approx 100)$, which sets ${\omega^{2}(\tilde{T})=\frac{3m^{2}M_{P}^{2}}{5\pi^{2}T_{c}^{4}\tilde{T}^{6}}}$. Initial conditions are set as : $\chi(\tilde{T}=1)=0$, $\chi^{\prime}(\tilde{T}=1)=-2$, for $\tilde{\phi}(\tilde{T}=1)=1$ and $\tilde{\phi^{\prime}}(\tilde{T}=1)=0$. The above equation further simplifies to a homogeneous one, as $\omega^{2}>>2$ for $\tilde{T}\leq 1$. Hence, the solution of $\chi$ is given by,
@@
Figure 1: Scenario-I: Evolution of the scalaron energy density (shown by the pink solid line) in the absence of non-minimal Higgs coupling to gravity. It first redshifts as radiation density (shown by the black dashed line), then as matter density (shown by the blue solid line). This feature is retained even with non-zero non-minimal Higgs coupling, as long as $3\xi m^{2}\neq\lambda v^{2}$.
@@
Noticeably, around $T\lesssim T_{c}~(=154~{\rm GeV})$, $\tilde{\phi}(\tilde{T})\sim\tilde{T}^{2}$, in turn $\rho_{\phi}\sim T^{4}$, while the second term in the equation being subdominant. This is a consequence of the scalaron-Higgs interaction, however small, keeps the scalaron coupled to the SM radiation bath for a significant duration of its evolution, depending on the scalaron mass in case of $\xi=0$. For $T<<T_{c}$, eventually the second term dominates, consequently the radiation-to-matter transition of the scalaron density takes place at temperature around $T_{d}=60$ MeV for $m=3.6\times 10^{-3}$ eV (obtained in Eq.24 to saturate present day DM density, discussed later), as shown in Fig.1. This also marks the decoupling of the scalaron from the Higgs, as the Higgs potential takes its usual $T=0$ form. Until the decoupling, the local minima of the scalaron shifts continuously and the scalaron tracks its time-dependent minima, with a characteristic oscillatory feature, apparent from Eq.22 and Eq.12. This is further depicted in Fig.2, by showing the evolution of $\phi$ (pink solid line) and its ambient minima (blue dashed line) at earlier epochs before the scalaron decoupling. We emphasize that the oscillatory feature shown in Fig.2 is inherent to scalaron dynamics since the initial epoch, unlike the standard axion scenarios.
Since, the scalaron energy density at late times, i.e. $T<<T_{c}$ matches with that of $\chi$, we can express the scalaron relic density in terms of the (scaled) energy density of $\chi$ at $T=T_{c}$ as
Figure 2: Scenario-I: Evolutions of the scalaron field (pink solid line) and its minima at early epochs. To note, the scalaron follows its minima that shifts continuously (depicted in blue dashed line) due to induced trilinear interaction with the SM Higgs.
@@
where $\rho_{c,0}(\approx 8.2\times 10^{-47}h^{-2}{~\rm GeV^{4}})$ and $T_{0}$ are the critical energy density and the photon temperature of the present universe, while $g_{*S}(T_{i})$ corresponds to the effective number of degrees of freedom contributing to the entropy density of the universe at $T=T_{i}$. In passing, we note that the decoupling temperature of the scalaron does not explicitly enter into the above expression; however, the decoupling must happen before the matter-radiation equality epoch, i.e. $T\sim 1$ eV, necessary condition for successful structure formation. In the present context, the decoupling happens around $60$ MeV, for which the scalaron density trivially satisfies the Big Bang Nucleosynthesis bound Fields:2019pfx on any extra radiation density present in the cosmic soup.
In calculating the relic density, the initial kinetic energy of $\chi$ only contributes, since at $T=T_{c}$, $\chi=0$, which renders $\rho_{\chi}(T_{c})=\dot{\chi}^{2}(T_{c})/2=2H^{2}_{c}\approx 2.05\times 10^{-27}~\rm{GeV^{2}}$. For $\xi=0$, $\phi_{0}(T_{c})=\frac{\lambda v^{4}}{2m^{2}M_{P}}=7.9\times 10^{-11}{\rm GeV}\left(\frac{{~\rm GeV}}{m}\right)^{2}$, and the observed DM relic density Planck:2018vyg is set by the scalaron mass only, i.e.
@@
where we have used $g_{*S}=3.91$ and $g_{*S}(T_{c})=g_{*}(T_{c})=100$. This reproduces the primary result of Refs.Shtanov:2021uif ; Shtanov:2024nmf ; Shtanov:2025nue within our general treatment of scalaron dynamics.
For non-zero values of $\xi$, particularly $\xi\ll\frac{\lambda v^{2}}{3m^{2}}$, the scalaron can in principle saturate the observed DM density only via the non-minimal Higgs coupling to gravity, as $\phi_{0}(T_{c})\approx 3\xi v^{2}/2M_{P}$. However, this would require an enormously large value of $\xi$ as shown in the following, which has been already excluded by the LHC constraints, as indicated earlier :
@@
Therefore, the scalaron dynamics for $3\xi m^{2}\neq\lambda v^{2}$, even with non-zero allowed values of $\xi$, is completely controlled by the Higgs quartic self-coupling. In fact, the initial energy density of the scalaron depends on $\phi_{0}$ and the Hubble parameter at $T=T_{c}$, which are determined by the SM value of $\lambda$. As a result, the scalaron mass is restricted to a value of $3.6\times 10^{-3}$ eV, to satisfy the DM relic density constraint. This restriction on the scalaron mass is lifted when the trilinear interaction becomes either sufficiently small or identically vanishes for $\lambda v^{2}=3\xi m^{2}$, which is studied next.
@@ 3.2 Scenario II : $3\xi m^{2}=\lambda v^{2}$
In this scenario, the trilinear interaction between Higgs and the scalaron can identically vanish for a wide range of allowed values of $\xi$, thereby the scalaron remains decoupled from the SM thermal bath at the outset of the evolution. Unlike the previous case, the scalaron density remains trivial if the scalaron stays at its minima, $\phi_{0}=0$. Therefore, the scalaron has to be misaligned from its minima with nearly zero velocity ($\dot{\phi}=0$), which facilitates coherent oscillation of the field around its trivial minima at a later time. This is apparent from the solution of Eq.17, i.e.
@@
In Fig.3, evolutions of the scalaron field (left panel) and its density (right panel) have been shown for a typical scalaron mass, $m=1$ keV. It is apparent that the oscillation starts around, $T_{osc}=4.96\times 10^{5}$ GeV, in agreement with Eq.27. As $\rho_{\phi}$ is approximately constant from $T_{i}=10^{7}$ GeV until $T=T_{osc}$, i.e. $\rho_{\phi}(T_{osc})\approx\rho_{\phi}(T_{i})$. So, the relic density can be expressed in terms of $\phi_{i}$ and $m$, utilizing Eq.27, as follows :
@@
where we have used $g_{*S}(T_{osc})=g_{*}(T_{osc})=100$. We can now choose $\xi$, which fixes the scalaron mass through the relation, $3\xi m^{2}=\lambda v^{2}$, since all other parameters are fixed at their SM value. Consequently, the remaining free parameter, $\phi_{i}$, is determined from the observed DM relic density. Unlike the previous scenario, the scalaron mass can vary depending on $\xi$, that can vary as high as $\mathcal{O}(10^{15})$. In the left panel of Fig.4, we depict the allowed parameter space of $m_{\phi}$ (and corresponding $\xi$ value) and $\phi_{i}$, for which the DM relic density is satisfied. In particular, the constant contour (blue solid line) of $\Omega_{DM}h^{2}=0.12$ has been shown depicting the central value of the DM relic density parameter. The region above the blue solid line in Fig.4 (left panel) is excluded, as the scalaron would be over-abundant, while the lower region corresponds to the scalaron contributing only a fraction of the DM abundance.
We note that the trilinear scalaron-higgs interaction vanishes only when the Higgs non-minimal coupling is introduced, thereby expanding the parameter space of the scalaron mass. Unlike scenario-I, here we have an extra freedom of choosing $\phi_{i}$ that opens a new parameter space for the scalaron mass to be probed in the near future.
Figure 4: Scalaron as dark matter in scenario-II : (Left panel) Constraints on the scalaron mass and initial field value from observed DM relic density, Higgs global signal strength and INTEGRAL/SPI limits on the excess photon flux from the scalaron decay to two photons. (Right panel) The scalaron decay rate to two photons and its exclusion limit at $95\%$ C.L., taken from Ref.Calore:2022pks . See text for details.
@@ 4 Constraints
In this section, we discuss both upper and lower limits of the scalaron mass for scenario-II, which are not solely determined from the observed DM relic density. The allowed scalaron mass, as we shall elucidate in the following subsections, varies from $1.38$ keV to $0.7$ MeV (see Fig.4). The lower limit of the scalaron mass is directly related to the upper bound on Higgs non-minimal coupling to gravity from LHC measurements via the relation $m_{\rm min}=(\lambda v^{2}/3\xi_{\rm max})^{1/2}$. Besides, the scalaron mediates an attractive Yukawa-like force between two non-relativistic particles, which is constrained from the torsion-balance experiment. The non-observation of the fifth force puts the lower limit on the scalaron mass as $2.7$ meV Kapner:2006si ; Adelberger:2006dh ; Cembranos:2008gj , which becomes redundant in the presence of LHC constraints on $\xi$ that sets $m_{\rm min}=1.38$ keV. The upper bound originates from the non-observation of an excess photon flux at INTEGRAL/SPI telescope, assuming the scalaron DM decaying to a pair of photons at the current epoch within our galaxy.
@@ 4.1 LHC Constraints on $\xi$
As already explained, the role of Higgs non-minimal coupling is pivotal in deciding the scalaron dynamics; however the bound on $\xi$ in the non-inflationary scenario is not well studied. Here we show the upper bound on $|\xi|$ from the measurement of Higgs global signal strength by LHC, following Ref.Atkins:2012yn . This bound is derived when the contribution from $R^{2}$-gravity becomes sub-dominant at $T\approx 0$ universe. Hence, the kinetic term of the SM Higgs is affected only by the non-minimal coupling, $\xi$, which requires the redefinition of the original Higgs field. As its result, Higgs couplings to all other SM particles are suppressed by a factor $\kappa$ Bezrukov:2007ep ; Atkins:2012yn given by,
$\kappa\approx\left(1+\frac{9\xi^{2}v^{2}}{M_{P}^{2}}\right)^{-1/2},$ (29)
while restoring canonical form of its kinetic term. The modification in Higgs couplings to other SM fields, in principle has imprints on the Higgs signal strength measurement at LHC, which is a standard practice to quantify any possible deviation from the SM prediction. In a narrow-width approximation of Higgs production cross-section and its decay, the Higgs signal strength of a process, $ii\rightarrow h\rightarrow ff$ is given by,
$\mu_{if}=\frac{\sigma(ii\rightarrow h)\cdot\mathrm{BR}(h\rightarrow ff)}{\sigma_{\mathrm{SM}}(ii\rightarrow h)\cdot\mathrm{BR}_{\mathrm{SM}}(h\rightarrow ff)}=\kappa^{2}\,$, (30)
where the deviation from the SM depends on $\kappa^{2}$ in our scenario. We use the global Higgs signal strength data, measured by both ATLAS ($\mu=1.05\pm 0.06$ ATLAS:2022vkf ) and CMS ( $\mu=1.002\pm 0.057$ CMS:2022dwd ) to estimate the upper bound on $\kappa^{2}$, and in turn on $|\xi|$. Combining these results, we find the upper limit on $|\xi|$ to be $1.25\times 10^{15}$ at the $95\%$ C.L., which results into the lower bound of the scalaron mass to be $1.38$ keV in scenario-II. This also explains why the scalaron density can not saturate the DM relic density in the scenario-I only with the Higgs non-minimal coupling, as shown in Eq.25.
4.2 Constraints on scalaron decay to photons

### v2

@@ 3 Scalaron dynamics : effects of Higgs interactions
With the general setup explained earlier, we now proceed to discuss two scenarios: one with a non-zero scalaron–Higgs interaction and the other with a vanishing scalaron–Higgs interaction. We solve Eq.16 and Eq.17 to demonstrate different dynamics at play, leading the scalaron to be a viable DM candidate. Consequently, the scalaron mass required to saturate the observed dark matter relic density differs widely, depending on the underlying dynamics.
@@ 3.1 Scenario I : $3\xi m^{2}\neq\lambda v^{2}$
In this scenario, the scalaron stays at its natural minima (as in Eq.12) before the EWPT and only starts moving as $T\leq T_{c}$. The scalaron, initially coupled to the SM thermal bath via the trilinear coupling, eventually decouples. Then it oscillates coherently as a matter field at temperatures well below $T_{c}$. This is a sheer contrast to axion misalignment scenario, in which the oscillation takes place depending only on the ratio of the Hubble parameter and the axion mass. Here, the onset of the oscillation is governed by the scalaron-Higgs interaction in addition to the Hubble to mass ratio. We demonstrate salient features of scalaron dynamics by solving Eq.16 semi-analytically with another change of variable, i.e., $\chi=\tilde{\phi}-\tilde{T}^{2}$, which yields,
@@
The equation can be solved analytically in the regime $\tilde{T}\leq 1$, for a constant $g_{*}(\approx 100)$, which sets ${\omega^{2}(\tilde{T})=\frac{3m^{2}M_{P}^{2}}{5\pi^{2}T_{c}^{4}\tilde{T}^{6}}}$. Initial conditions are set as : $\chi(\tilde{T}=1)=0$, $\chi^{\prime}(\tilde{T}=1)=-2$, for $\tilde{\phi}(\tilde{T}=1)=1$ and $\tilde{\phi^{\prime}}(\tilde{T}=1)=0$. The above equation further simplifies to a homogeneous one, as $\omega^{2}\gg 2$ for $\tilde{T}\leq 1$. Hence, the solution of $\chi$ is given by,
@@
Figure 1: Scenario-I: Evolution of the scalaron energy density (shown by the pink solid line) in the absence of non-minimal Higgs coupling to gravity. It first redshifts as radiation density (shown by the black dashed line), then as matter density (shown by the blue solid line). This feature is retained even with non-zero non-minimal Higgs coupling, as long as $3\xi m^{2}\neq\lambda v^{2}$. See text for details.
@@
Noticeably, around $T\lesssim T_{c}~(=154~{\rm GeV})$, $\tilde{\phi}(\tilde{T})\sim\tilde{T}^{2}$, in turn $\rho_{\phi}\sim T^{4}$, while the second term in the equation being subdominant. This is a consequence of the scalaron-Higgs interaction, however small, keeps the scalaron coupled to the SM radiation bath for a significant duration of its evolution, depending on the scalaron mass in case of $\xi=0$. For $T\ll T_{c}$, eventually the second term dominates, consequently the radiation-to-matter transition of the scalaron density takes place at temperature around $T_{d}=60$ MeV for $m=3.6$ meV (obtained in Eq.24 to saturate present day DM density, discussed later), as shown in Fig.1. This also marks the decoupling of the scalaron from the Higgs, as the Higgs potential takes its usual $T=0$ form. Until the decoupling, the local minima of the scalaron shifts continuously and the scalaron tracks its time-dependent minima, with a characteristic oscillatory feature, apparent from Eq.22 and Eq.12. This is further depicted in Fig.2, by showing the evolution of $\phi$ (pink solid line) and its ambient minima (blue dashed line) at earlier epochs before the scalaron decoupling. We emphasize that the oscillatory feature shown in Fig.2 is inherent to scalaron dynamics since the initial epoch, unlike the standard axion scenarios. In passing, we also note that these features of scalaron is retained in presence of non-zero allowed values of $\xi$, as long as the trilinear interaction with the Higgs is non-trivial.
Since, the scalaron energy density at late times, i.e., $T\ll T_{c}$ matches with that of $\chi$, we can express the scalaron relic density in terms of the (scaled) energy density of $\chi$ at $T=T_{c}$ as
Figure 2: Scenario-I: Evolutions of the scalaron field (pink solid line) and its minima (blue dashed line) at early epochs. To note, the scalaron follows its minima that shifts continuously due to induced trilinear interaction with the SM Higgs. See text for details.
@@
where $\rho_{c,0}(\approx 8.2\times 10^{-47}h^{-2}{~\rm GeV^{4}})$ and $T_{0}$ are the critical energy density and the photon temperature of the present universe, while $g_{*S}(T_{i})$ corresponds to the effective number of degrees of freedom contributing to the entropy density of the universe at $T=T_{i}$. In passing, we note that the decoupling temperature of the scalaron does not explicitly enter into the above expression; however, the decoupling must happen before the matter-radiation equality epoch, i.e., $T\sim 1$ eV, necessary condition for successful structure formation. In the present context, the decoupling happens around $60$ MeV (with $\xi=0$ and $m=3.6$ meV), for which the scalaron density trivially satisfies the Big Bang Nucleosynthesis bound Fields:2019pfx on any extra radiation density present in the cosmic soup.
In calculating the relic density, the initial kinetic energy of $\chi$ only contributes, since at $T=T_{c}$, $\chi=0$, which renders $\rho_{\chi}(T_{c})=\dot{\chi}^{2}(T_{c})/2=2H^{2}_{c}\approx 2.05\times 10^{-27}~\rm{GeV^{2}}$. In addition, we need the initial field value, $\phi(T_{c})$, which is determined by both $\lambda$ and $\xi$. Depending on their relative contribution, there are two distinct subclasses within scenario-I, namely $\xi\ll\frac{\lambda v^{2}}{3m^{2}}$ and $\xi\gg\frac{\lambda v^{2}}{3m^{2}}$.
For $\xi\ll\frac{\lambda v^{2}}{3m^{2}}$, $\phi_{0}(T_{c})=\frac{\lambda v^{4}}{2m^{2}M_{P}}=7.9\times 10^{-11}{\rm GeV}\left(\frac{{~\rm GeV}}{m}\right)^{2}$, and the observed DM relic density Planck:2018vyg is set by the scalaron mass only, i.e.,
@@
where we have used $g_{*S}=3.91$ and $g_{*S}(T_{c})=g_{*}(T_{c})=100$. This reproduces the primary result of Refs.Shtanov:2021uif ; Shtanov:2024nmf ; Shtanov:2025nue with $\xi=0$ within our general treatment of scalaron dynamics.
For $\xi\gg\frac{\lambda v^{2}}{3m^{2}}$, the scalaron starts from a negative field value, $\phi_{0}(T_{c})\approx-3\xi v^{2}/2M_{P}$, unlike the previous one. In this case, the observed DM relic density is saturated by $\xi$ only, as follows :
@@
Is such a large required value of $\xi$ allowed in the scalaron-Higgs mixed model ? We have discussed the issue in great detail in Sec.4.1 and its implication on the scalaron parameter space in Sec. 5. We shall see that indeed such a large value of $\xi$ is allowed as long as the scalaron mass is sufficiently small.
Noticeably, the scalaron dynamics is completely controlled by the Higgs couplings ($\lambda$ and $\xi$) for $3\xi m^{2}\neq\lambda v^{2}$. In fact, the initial energy density of the scalaron depends on the Hubble parameter at $T=T_{c}$ and $\phi_{0}$ which is determined by the non-minimal coupling $\xi$ and the SM value of $\lambda$. The scalaron saturates the DM density for a relatively wide range of masses when the trilinear interaction becomes either sufficiently small or vanishes (to a good approximation) for $\lambda v^{2}=3\xi m^{2}$, which is studied next.
To note, the scenario-I is studied assuming the EWPT as a second order transition, which is a valid assumption for scalaron mass around a few meVShtanov:2021uif ; Shtanov:2022xew . However, in Ref.Shtanov:2024nmf , the author points out how the scalaron dynamics needs to be modified for higher masses with a realistic scenario of the EWPT, which is a smooth cross-over. In addition, effects due to the QCD trace anomaly has also been discussed in Ref.Shtanov:2024nmf .
@@ 3.2 Scenario II : $3\xi m^{2}=\lambda v^{2}$
In this scenario, the trilinear interaction between Higgs and the scalaron can vanish at the leading order for a wide range of allowed values of $\xi$ and $m$, thereby the scalaron remains decoupled from the SM thermal bath at the outset of the evolution. Hence, its dynamics is not controlled by the EWPT, unlike scenario-I. The scalaron density remains trivial if it stays at its minima, $\phi_{0}=0$. Therefore, the scalaron has to be misaligned from its minima with nearly zero velocity ($\dot{\phi}=0$), which facilitates coherent oscillation of the field around its trivial minima at a later time. This is apparent from the solution of Eq.17, i.e.,
@@
In Fig.3, evolutions of the scalaron field (left panel) and its density (right panel) have been shown for a typical scalaron mass, $m=1$ keV. It is apparent that the oscillation starts around, $T_{osc}=4.96\times 10^{5}$ GeV, in agreement with Eq.27. As $\rho_{\phi}$ is approximately constant from $T_{i}=10^{7}$ GeV until $T=T_{osc}$, i.e., $\rho_{\phi}(T_{osc})\approx\rho_{\phi}(T_{i})$. So, the relic density can be expressed in terms of $\phi_{i}$ and $m$, utilizing Eq.27, as follows :
@@
where we have used $g_{*S}(T_{osc})=g_{*}(T_{osc})=100$. We can now choose $\xi$, which fixes the scalaron mass through the relation, $3\xi m^{2}=\lambda v^{2}$, since all other parameters are fixed at their SM value. Consequently, the remaining free parameter, $\phi_{i}$, is determined from the observed DM relic density. In this case, the scalaron mass can vary in a wide range of values with corresponding value of $\phi_{i}$, unlike the previous scenario, especially for $\xi\ll\lambda v^{2}/3m^{2}$. We demonstrate the allowed parameter space of $m$ and $\phi_{i}$ in Sec.5 and in Fig.6.
@@ 4 Constraints
Having explained the scalaron dynamics in the previous section, we now elucidate the generic constraints on the scalaron mass $m$ and the Higgs non-minimal coupling to gravity $\xi$, and summarize the results using Fig.5. These constraints are common to both scalaron dynamics scenarios considered in this work. We provide a detailed discussion on these constraints in the subsequent subsections 4.1, 4.2 and 4.3.
@@ 4.1 LHC Constraints on $\xi$
As already explained, the role of Higgs non-minimal coupling is pivotal in deciding the scalaron dynamics; however the bound on $\xi$ in the non-inflationary scenario is not well studied. In Ref.Atkins:2012yn the upper bound on $|\xi|$ derived from the Higgs global signal strength measurement by the LHC has been reported, in the absence of $R^{2}$-gravity. In the current situation, the scalaron entangles both $R^{2}$-gravity contribution and the contribution from Higgs non-minimal coupling to gravity, for which an independent bound on $|\xi|$ becomes obscure. However, we find the upper bound on $|\xi m|$ from Higgs measurements, after making appropriate field transformations, details of which are given below.
We first write the conformal factor of Eq.4 introducing a new field, $\psi=-\frac{\sigma M_{p}}{3m^{2}}$, as the following
$\Omega^{2}=\left[1+\frac{\psi}{M_{P}}+\frac{3\xi}{M^{2}_{P}}\left(\Phi^{\dagger}\Phi-\frac{v^{2}}{2}\right)\right],$ (29)
where $\psi$ represents the scalar degree of freedom associated with $R^{2}$-gravity. With this, we can recast the matter part of Eq.5 in terms of $\psi$ and the bare Higgs field as the following.
$S_{M}=\frac{M^{2}_{P}}{3}\int\frac{d^{4}x\sqrt{-\tilde{g}}}{\Omega^{4}}\Bigg[$ $\Omega^{2}\tilde{g}^{\mu\nu}\left(D_{\mu}\Phi\right)^{\dagger}D_{\nu}\Phi+\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}\psi\partial_{\nu}\psi+\frac{9\xi^{2}}{2M_{P}^{2}}\tilde{g}^{\mu\nu}\partial_{\mu}\left(\Phi^{\dagger}\Phi\right)\partial_{\nu}\left(\Phi^{\dagger}\Phi\right)$
$+\frac{3\xi}{M_{P}}\tilde{g}^{\mu\nu}\partial_{\mu}\psi\partial_{\nu}\left(\Phi^{\dagger}\Phi\right)-\frac{1}{2}m^{2}\psi^{2}-\lambda\left({\Phi}^{\dagger}{\Phi}-\frac{v^{2}}{2}\right)^{2}\bigg].$ (30)
We note that a kinetic mixing between $\psi$ and the Higgs emerges at $T\leq T_{c}$, via the non-minimal Higgs coupling, $\xi$. Our main interest is in the quadratic part of the theory, which is given by,
$\mathcal{L}_{2}=\frac{1}{2}\left(1+{9x^{2}}\right)\partial_{\mu}\zeta\partial^{\mu}\zeta+\frac{1}{2}\partial_{\mu}\psi\partial^{\mu}\psi+{3x}\partial_{\mu}\zeta\partial^{\mu}\psi-\frac{1}{2}\left(M_{\zeta}^{2}\zeta^{2}+m^{2}\psi^{2}\right),$ (31)
where we have used, $\Phi^{T}=(0~~(v+\zeta)/\sqrt{2})$, $x=\xi v/M_{P}$ and $M_{\zeta}=\sqrt{2\lambda}v$. Evidently, the redefinition of the kinetic term of the Higgs can not be performed independently as in Refs.Bezrukov:2007ep ; Atkins:2012yn ; Boyanovsky:2017esz due to the kinetic mixing. Therefore, we make the following basis transformation to get canonical kinetic terms for both the fields.
$\eta_{1}$ $=\sqrt{1+9x^{2}}\left(\zeta+\frac{3x}{1+9x^{2}}\psi\right),$ (32)
$\eta_{2}$ $=\frac{\psi}{\sqrt{1+9x^{2}}}.$
Although we get rid of the kinetic mixing, mass mixing between two fields appear in the new basis, as follows :
$\mathcal{L}_{2}=\frac{1}{2}\partial_{\mu}\eta_{1}\partial^{\mu}\eta_{1}+\frac{1}{2}\partial_{\mu}\eta_{2}\partial^{\mu}\eta_{2}-\frac{1}{2}M_{1}^{2}\eta_{1}^{2}-\frac{1}{2}M_{2}^{2}\eta_{2}^{2}-M_{12}^{2}\eta_{1}\eta_{2},$ (33)
where,
$M_{1}^{2}$ $=\frac{M_{\zeta}^{2}}{1+9x^{2}},$ (34)
$M_{2}^{2}$ $=\frac{9x^{2}M_{\zeta}^{2}}{1+9x^{2}}+m^{2}\left(1+9x^{2}\right),$
$M_{12}^{2}$ $=-\frac{3xM_{\zeta}^{2}}{1+9x^{2}}.$
Now, we introduce a rotation of the fields to diagonalize the mass matrix, $M_{S}$, i.e.,
[…]
where the mixing angle and the mass eigenvalues are given by,
[…]
$\mathcal{M}^{2}_{1,2}=\frac{1}{2}\left[M_{\zeta}^{2}+m^{2}\left(1+9x^{2}\right)\pm\sqrt{M_{\zeta}^{4}+m^{4}\left(1+9x^{2}\right)^{2}+2m^{2}M_{\zeta}^{2}\left(9x^{2}-1\right)}\right].$ (36)
With these transformations, we arrive at a basis in which the kinetic terms are canonical and the mass matrix is diagonal. We identify $\phi_{1}$ as the SM Higgs-like field with mass eigenvalue
$\mathcal{M}_{1}=M_{\zeta}\sqrt{1+\frac{9x^{2}m^{2}}{M_{\zeta}^{2}}},$ (37)
where we have assumed $m\ll M_{\zeta}$. This assumption is consistent with the scalaron dynamics and, in particular, prevents the decay of the scalaron into two Higgs particles. Up to this point, we have not imposed any assumption on $x$ and have kept it arbitrary. However, for $\phi_{1}$ to be the physical Higgs and being consistent with the scalaron dynamics explained in previous sections, $|xm/M_{\zeta}|$ should be small. This can be readily seen from the expressions of the mixing angle in Eq. (36) approximated as,
[…]
Using these relations, the new fields can be expressed in terms of the original ones as
[…]
To note, the cosmology of $\phi_{2}$ field will be similar to original scalaron field ($\phi$), as defined in Sec.2, since the energy density is determined by its mass, which is identical to the scalaron mass, i.e., $\mathcal{M}_{2}=m$. This is also apparent from the above relations that $\phi_{1}$ acts just like the SM Higgs and $\phi_{2}$ as scalaron, having contribution from both Higgs and the additional scalar degree from $R^{2}$ gravity.
The advantage of this new basis is that we get our SM-like Higgs in its mass (physical) basis with canonical kinetic terms, which is essential to derive relevant bounds from experimental results on new physics scenarios, like the present one. In the present theory, the original Higgs field is just rotated by a factor $\kappa$ (shown in Eq.41), since $3xm^{2}/M^{2}_{\zeta}\rightarrow 0$, even with highest possible scalaron mass, i.e., $m=0.7$ MeV, stemming independently from the indirect detection constraints. The effect is captured in the Higgs signal strength measurement at LHC, which is a standard practice to quantify any possible deviation from the SM prediction. In a narrow-width approximation of Higgs production cross-section and its decay, the Higgs signal strength of a process, $ii\rightarrow h\rightarrow ff$ is given by,
$\mu_{if}=\frac{\sigma(ii\rightarrow h)\cdot\mathrm{BR}(h\rightarrow ff)}{\sigma_{\mathrm{SM}}(ii\rightarrow h)\cdot\mathrm{BR}_{\mathrm{SM}}(h\rightarrow ff)}.$ (40)
In our scenario, $\mu_{if}=\kappa^{2}$ and it is given by,
$\kappa^{2}=\left(1+\frac{18m^{2}x^{2}}{M_{\zeta}^{2}}\right).$ (41)
Using the Higgs mass measurement at the LHC, $\mathcal{M}_{1}=125.11\pm 0.11$ GeV ATLAS:2023oaq , obtained from the combined analysis of the decay channels $H\to ZZ^{*}\to 4\ell$ and $H\to\gamma\gamma$, together with the global Higgs signal strength measurements by ATLAS ($\mu=1.05\pm 0.06$ ATLAS:2022vkf ) and CMS ($\mu=1.002\pm 0.057$ CMS:2022dwd ), we can obtain allowed region of $\lambda$ and $|xm|$, which are free parameters of the theory in its current form. However, we have checked numerically that $\lambda$ is restricted within a very small region, i.e. $0.12\lesssim\lambda\lesssim 0.13$ to satisfy the Higgs mass and the global signal strength simultaneously. Therefore, we obtain the bound on $|xm|$ to be $2.5$ GeV taking $\lambda=0.13$, which is also used in previous sections. This bound can vary only up to $10$ GeV for $\lambda=0.12$. Therefore, we find the bound on $|\xi m|$ in the scalaron-Higgs mixed model from LHC measurement as, $|\xi m|<2.35\times 10^{16}~\mathrm{GeV}$. Implications of this bound on scalaron DM scenarios have been discussed in Sec.5 and in Fig.5. We note that the non-minimal coupling $\xi$ can take very large values, exceeding the bound reported in Ref. Atkins:2012yn by several orders of magnitude, provided the scalaron mass is sufficiently small. Furthermore, the perturbativity of the scalaron-Higgs model (particularly for Higgs quartic self coupling Ema:2017rqn ) is also ensured with the current bound, i.e., $\xi^{2}m^{2}/M^{2}_{P}\approx 10^{-4}\ll 4\pi$.
Figure 4: Constraint on the scalaron decay rate : The scalaron decay rate (red solid line) as a function of its mass, compared with the exclusion limit at $95\%$ C.L. (green shaded region) on the light DM decaying into a pair of photons, taken from Ref.Calore:2022pks . See text for details.
4.2 Constraints on the scalaron decay to photons